# Shangjiong Wang1, Zhihong Man1, Zhengwei Cao1, Jinchuan Zheng1 and Zhikang Ge2 1School of Science, Computing and Engineering Technologies

### v1

@@ Abstract
Multivariate time series forecasting relies on accurately capturing the correlations among variates. Current channel-independent (CI) models and models with a CI final projection layer are unable to capture these dependencies. In this paper, we present the variate embedding (VE) pipeline, which learns a unique and consistent embedding for each variate and combines it with Mixture of Experts (MoE) and Low-Rank Adaptation (LoRA) techniques to enhance forecasting performance while controlling parameter size. The VE pipeline can be integrated into any model with a CI final projection layer to improve multivariate forecasting. The learned VE effectively groups variates with similar temporal patterns and separates those with low correlations. The effectiveness of the VE pipeline is demonstrated through extensive experiments on four widely-used datasets. The code is available at: https://github.com/swang-song/VE.
@@ I Introduction
After obtaining the temporal projection embedding, there are channel-independent (CI) models that treat all variates the same [2, 3] and attention models that leverage the attention mechanism [4] to exploit multivariate dependencies [1, 5]. However, both methods have limitations. CI models, due to weight sharing among all variates, can only learn general patterns and cannot model patterns unique to specific variates as the distinct pattern in one variate is a noise for other variates with low correlation. Meanwhile, attention models fail to provide a unified and consistent representation for each variate, as the attention maps evolve across layers and become more abstract in deeper layers. Additionally, even with the attention mechanism, the final projection layer remains channel-independent, which is unable to react to variate-specific patterns. Therefore, we pose the question: Can we train a unique and consistent representation for each variate and utilize this in CI models or layers to improve forecasting?
@@
  • We propose variate embedding (VE), a new paradigm for multivariate time series modeling, in addition to existing temporal projection embedding and channel projection embedding methods.
@@
  • We conduct extensive evaluations of VE on four commonly used datasets, demonstrating its effectiveness in modeling multivariate correlations and improving forecasting while maintaining parameter efficiency.
@@ II Related Works
Mixture-of-Expert (MOE) layers consist of multiple sub-networks and each input is processed by an input-dependent combination of these experts [6]. MoE has been utilized in Fedformer [17] for adaptively selecting the pooling filters to extract trand from time series. Moreover, MoLE [18] applied MoE to the DLinear [2] and the experts are time-stamp dependent.
@@ III-A VEMoE
$\hat{\bm{Y}}=\{\tilde{\bm{W}}_{1}\bm{x}_{1},\dots,\tilde{\bm{W}}_{C}\bm{x}_{C}\}$ (2)
@@ III-B LoRA
TABLE I: Comparison of baselines and proposed VE on ETTh1, ETTh2, ECL and Weather with prediction horizon $H\in\{92,192\}$. The best results are highlighted in bold. The std values are listed in Table V
@@ IV-A Experimental Setup
Datasets. We evaluate our VE pipeline on four popular and publicly available real-world datasets: ETTh1, ETTh2, ECL and Weather [21]. Detailed descriptions are in Appendix -B
@@
  • DLinear [2]: This is used to demonstrate the case where there are multiple inputs with different projection weights in the last layer and the prediction is the sum of these projections. We treat these multiple weights as a group and train a single VE for all the weights within the same group, as as using separate VEs for different weights in a group did not yield statistically significant improvements and only increased the parameter count in our experiments.
@@ IV-B Main Results
TABLE II: Comparison of baselines and proposed VE in mixed dataset with prediction horizon $H\in\{92,192\}$ on mixed dataset. The best results are highlighted in bold.
@@
Last layer linear VE linear VE
@@ IV-D Abalation study
The default FITS model has 48.86K parameters and an MSE of 0.331. Our VEMoE alone improves forecasting but increases parameter size significantly. Adding LoRA reduces both the MSE and parameter count. Specifically, without LoRA, the improvement plateaus at an MSE of 0.328 with 16 times ($k=16$) the original FITS parameters. With LoRA, the MSE remains at 0.328, but with only 43K parameters ($p=1$). Remarkably, using just 10K parameters ($p=0.25$), the MSE of 0.332 is still competitive.
@@ V Conclusion
This paper presents the VE pipeline, designed to capture multivariate correlations through the proposed variate embedding. By combining VE with a MoE architecture, the model can learn optimal weights for each variate, while low-rank adaptation (LoRA) is employed to improve parameter efficiency. Our VE pipeline is applicable to any model with a channel-independent final projection layer, addressing the limitation of such layers in learning unique patterns specific to each variate. The effectiveness and compatibility of the proposed VE pipeline are validated through extensive experiments on four widely-used datasets. We believe that our VE approach has the potential to be beneficial in other time series analysis tasks, such as time series imputation and anomaly detection, and building parameter-efficient time series foundation models to process various time series tasks and diverse data domains.

### v2

# Shangjiong Wang1, Zhihong Man1, Zhenwei Cao1, Jinchuan Zheng1 and Zhikang Ge2 1School of Science, Computing and Engineering Technologies

@@ Abstract
Multivariate time series forecasting relies on accurately capturing the correlations among variates. Current channel-independent (CI) models and models with a CI final projection layer are unable to capture these dependencies. In this paper, we present the variate embedding (VE) pipeline, which learns a unique and consistent embedding for each variate and combines it with Mixture of Experts (MoE) and Low-Rank Adaptation (LoRA) techniques to enhance forecasting performance while controlling parameter size. The VE pipeline can be integrated into any model with a CI final projection layer to improve multivariate forecasting. The learned VE effectively groups variates with similar temporal patterns and separates those with low correlations. The effectiveness of the VE pipeline is demonstrated through experiments on four widely-used datasets. The code is available at: https://github.com/swang-song/VE.
@@ I Introduction
After obtaining the temporal projection embedding, there are channel-independent (CI) models that treat all variates the same [2, 3] and attention models that leverage the attention mechanism [4] to exploit multivariate dependencies [1, 5]. However, both methods have limitations. CI models, due to weight sharing among all variates, can only learn general patterns and cannot model patterns unique to specific variates as the distinct pattern in one variate is a noise for other variates with low correlation. Meanwhile, attention models fail to provide a unified and consistent representation for each variate, as the attention maps evolve across heads and layers and become more abstract in deeper layers. Additionally, even with the attention mechanism, the final projection layer remains channel-independent, which is unable to react to variate-specific patterns. Therefore, we pose the question: Can we train a unique and consistent representation for each variate and utilize this in CI models or layers to improve forecasting?
@@
  • We propose variate embedding (VE), a new paradigm for multivariate time series modeling, on top of the existing temporal projection embedding and channel projection embedding methods.
@@
  • We conduct experiments of VE on four commonly used datasets, demonstrating its effectiveness in modeling multivariate correlations and improving forecasting while maintaining parameter efficiency.
@@ II Related Works
Mixture-of-Expert (MOE) layers consist of multiple sub-networks and each input is processed by an input-dependent combination of these experts [6]. MoE has been utilized in Fedformer [17] for adaptively selecting the pooling filters to extract trend from time series. Moreover, MoLE [18] applied MoE to the DLinear [2] and the experts are time-stamp dependent.
@@ III-A VEMoE
$\bm{Y}=\{\tilde{\bm{W}}_{1}\bm{x}_{1},\dots,\tilde{\bm{W}}_{C}\bm{x}_{C}\}$ (2)
@@ III-B LoRA
TABLE I: Comparison of baselines and proposed VE on ETTh1, ETTh2, ECL and Weather with prediction horizon $H\in\{92,192\}$. The best results are highlighted in bold.
@@ IV-A Experimental Setup
Datasets. We evaluate our VE pipeline on four popular and publicly available real-world datasets: ETTh1, ETTh2, ECL and Weather [21].
@@
  • DLinear [2]: This is used to demonstrate the case where there are multiple inputs with different projection weights in the last layer and the prediction is the sum of these projections. We treat these multiple weights as a group and train a single VE for all the weights within the same group, as as using separate VEs for different weights in a group did not yield statistically significant improvements and only increased the parameter size in our experiments.
@@ IV-B Main Results
TABLE II: Comparison of baselines and proposed VE on mixed dataset with prediction horizon $H\in\{92,192\}$. The best results are highlighted in bold.
@@
Last layer original VE original VE
@@ IV-D Abalation study
The default FITS model has 48.86K parameters and an MSE of 0.331. Our VEMoE alone improves forecasting but increases parameter size significantly. Adding LoRA reduces both the MSE and parameter size. Specifically, without LoRA, the improvement plateaus at an MSE of 0.328 with 16 times ($k=16$) the original FITS parameters. With LoRA, the MSE remains at 0.328, but with only 43K parameters ($p=1$). Remarkably, using just 10K parameters ($p=0.25$), the MSE of 0.332 is still competitive.
@@ V Conclusion
This paper presents the VE pipeline, designed to capture multivariate correlations through the proposed variate embedding. By combining VE with a MoE architecture, the model can learn optimal weights for each variate, while low-rank adaptation (LoRA) is employed to improve parameter efficiency. Our VE pipeline is applicable to any model with a channel-independent final projection layer, addressing the limitation of such layers in learning unique patterns specific to each variate. The effectiveness and compatibility of the proposed VE pipeline are validated through experiments on four widely-used datasets. We believe that our VE approach has the potential to be beneficial in other time series analysis tasks, such as time series imputation and anomaly detection, and building parameter-efficient time series foundation models to process various time series tasks and diverse data domains.